OpenStudy (anonymous):

2x-6=56-x^2 @ 2x-6=56-x^2 @ @Mathematics

OpenStudy (anonymous):

(x2+2x+3)(x2+4x+3)>0

OpenStudy (saifoo.khan):

2x-6=56-x^2 x^2 + 2x - 6 - 56 = 0 x^2 + 2x - 72 = 0

OpenStudy (saifoo.khan):

What do u want to find?

OpenStudy (waheguru):

http://openstudy.com/#/updates/4ec18e80e4b032a1a59a5774 help

OpenStudy (anonymous):

x

OpenStudy (anonymous):

determine

jhonyy9 (jhonyy9):

why 72 ? there is 62

jhonyy9 (jhonyy9):

-62